OpenStudy (anonymous):

Find the difference 7x - (25x^2 + 12x) =

OpenStudy (cwrw238):

first you need to distribute -1 over the contents of the parentheses

OpenStudy (anonymous):

30x^2 -25x^2-5x -25x^2+19x 25x^2-5x

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

so what happens to the +12x?

OpenStudy (cwrw238):

when you multiply it by -1 ?

OpenStudy (anonymous):

It becomes negative i guess

OpenStudy (cwrw238):

right -12x so you have 7x - 25x^2 - 12x now simplify this

OpenStudy (cwrw238):

and you have the solution

OpenStudy (anonymous):

Would that be c??

OpenStudy (cwrw238):

no you need to work out 7x - 12x

OpenStudy (cwrw238):

what does that come to?

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

7x - (25x^2 + 12x) = 7x - 25x^2 - 12x = -25x^2 - 5x = -25x^2 - 5x

OpenStudy (anonymous):

Thanks

OpenStudy (cwrw238):

can you understand why 7x - 12x = -5x?

OpenStudy (anonymous):

Yea haha i accidentally added instead of subtracted

OpenStudy (cwrw238):

ok